OpenStudy (anonymous):

The parabola y=ax^2+bx+c, c≠0 meets the x-axis at A(α,0) and B(β,0) and the y-axis at C. If AC and BC are perpendicular, prove ac=-1

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

alpha and beta are the roots of quadratic ax^2+bx+c=0

OpenStudy (anonymous):

so alpha + beta = -b/a and alpha* beta= c/a

OpenStudy (lgbasallote):

sure...tag the math hater to a math problem...makes a lot of sense

OpenStudy (anonymous):

slope of AC * slope of BC= -1

OpenStudy (anonymous):

so alpha* beta = -c^2

OpenStudy (anonymous):

and put this value into... alpha* beta= c/a

OpenStudy (anonymous):

so ac=-1

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

how did you get to alpha*beta=-c^2?

OpenStudy (anonymous):

use this...slope of AC * slope of BC= -1 c=(0, beta)

mathslover (mathslover):

No one tags mathslover though :P

OpenStudy (anonymous):

since when does c=(0,beta)?

OpenStudy (anonymous):

so slope of AC= - c/alpha=m1 , slope of BC= -c/beta=m2 m1 *m2= -1...since AC is perpendicular to BC.

OpenStudy (anonymous):

sorry C=(0,c)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

@akash123 @ganeshie8

OpenStudy (anonymous):

u got it?

OpenStudy (anonymous):

yah